An effective rate of 29.61% corresponds to a nominal rate, compounded weekly,of?

*34.35%

*29.53%

*26%

*29.61%






1
Expert's answer
2022-02-28T11:41:04-0500

"i=100((\frac{1+j}{m})^m-1)"


"29.61=100((1+\frac {r} {52} )^{52}-1 )"


r=26%